OpenStudy (anonymous):

Graph the six terms of a finite series where a1 = -3 and r = 1.5.

OpenStudy (anonymous):

graphed sequence showing point 1, 2.25, point 2, 3.75, point 3, 8.25, point 4, 21.75, point 5, 62.25, and point 6, 183.75 graphed sequence showing point 1, 1.5, point 2, 3.75, point 3, 10.5, point 4, 30.75, point 5, 91.5, and point 6, 273.75 graphed sequence showing point 1, negative 9, point 2, negative 12, point 3, negative 16.5, point 4, negative 23.25, point 5, negative 33.38, and point 6, negative 48.56 graphed sequence showing point 1, negative 3, point 2, negative 7.5, point 3, negative 14.25, point 4, negative 24.38, point 5, negative 39.56, and point 6, negative 62.34

OpenStudy (anonymous):

@johnweldon1993 can u plz help me out?!

OpenStudy (triciaal):

one approach this is a geometric progression the nth term a^n = a1*r*(n-1) a1 is the 1st term r is the common ratio between the terms to get the 6 terms find a when n = 1 then when n = 2 etc a^n = y and n = x graph the points (x, y )

OpenStudy (anonymous):

@e.mccormick